In 1 L of Mannitol Salt Agar there is 1 g of beef extract. Convert this to a weight over volume percentage (%w/v) and report your answer rounded to two decimal places. Do not include units in the answer block. Percent weight over volume (%w/v) is the weight of the solute in grams divided by the volume of the solvent in mL times 100. It is essentially expressing the grams of solute per 100 mL of solvent.

In 1 L of Eosine Methylene Blue Agar there are 0.065 g of methylene blue. Convert this to a weight over volume percentage (%w/v) and report your answer rounded to FOUR decimal places. Do not include units in the answer block.

Percent weight over volume (%w/v) is the weight of the solute in grams divided by the volume of the solvent in mL times 100. It is essentially expressing the grams of solute per 100 mL of solvent.

In 1 L of MacConkey Agar there are 10 g of lactose. Convert this to a weight over volume percentage (%w/v) and report your answer rounded to two decimal places. Do not include units in the answer block.

Percent weight over volume (%w/v) is the weight of the solute in grams divided by the volume of the solvent in mL times 100. It is essentially expressing the grams of solute per 100 mL of solvent.
Answer #1

1L of Mannitol Salt Agar has 1g Beef Extract.

% w/v = (1 * 100)/ 1000 = 0.10 %

1 L of Eosine Methylene Blue Agar has 0.065 g of methylene blue.

% w/v = (0.065 * 100)/1000 = 0.0065 %

1 L of MacConkey Agar has 10 g of lactose.

% w/v = (10 * 100)/1000 = 1.00 %